# Equilibrium under Time-Inconsistency: A New Existence Theory by Vanishing Entropy Regularization

### 11 Mar 2026 v1

@@ 2 Model Setup
We consider a time-inconsistent stochastic control problem. Let $U$ denote the action space, which is a bounded subset of $\mathbb{R}^{\ell}$ with $\operatorname{\texttt{Leb}}(U)>0$. We also denote by ${\mathcal{P}}(U)$ as the set of all probability measures on $U$ and ${\mathcal{P}}_{c}(U)$ as the subset of ${\mathcal{P}}(U)$ which contains all probability measures absolutely continuous with respect to (w.r.t.) the Lebesgue measure. For technical convenience, we focus on the model over an infinite horizon where only the drift coefficient of the state process is controlled. The general case when the diffusion coefficient is also controlled will be left for future study.
@@
admits a unique strong solution, where $b:\mathbb{R}^{d}\times U\mapsto\mathbb{R}^{d}$, and $\sigma:\mathbb{R}^{d}\mapsto\mathbb{R}^{d}\times\mathbb{R}^{m}$ and $W$ is a standard $m$-dimensional Brownian motion. The value function under the relaxed control $\pi$ over an infinite horizon is defined by
$V^{\pi}(t,x)=\mathbb{E}_{x}\left[\int_{0}^{\infty}r^{\pi}(t+s,X^{\pi}_{s})ds\right],$
where $r(t,x,a):\mathbb{T}\times\mathbb{R}^{d}\times U\mapsto\mathbb{R}$ is the reward function. Let $J^{\pi}(x):=V^{\pi}(0,x)$ be the value function at the initial time $t=0$ under the relaxed control $\pi$. Given generic functions $f(t,x,a):\mathbb{T}\times{\mathbb{R}^{d}}\times U\mapsto\mathbb{R}$ and $\varpi\in{\mathcal{P}}(U)$, let us define
$f^{\varpi}(t,x):=\int_{U}f(t,x,a)\varpi(da).$
@@
$K_{2}:=\int_{0}^{\infty}\sup_{x\in\mathbb{R}^{d},a\in U,u\geq s}\frac{|\partial_{t}r(s,x,a)-\partial_{t}r(u,x,a)|}{|s-u|^{\alpha/2}}ds<\infty.$ (2.2)
@@
The first boundedness condition in (2.1) ensures that $J^{\pi}$ is well-defined for any admissible relaxed control $\pi$, (2.2) is easily satisfied if the second derivative of $r$ on $t$ is integrable over the time horizon $\mathbb{T}$ uniformly over all $a,x$. This assumption will be used later to deliver the Hölder continuity of $\partial_{t}V^{\pi}$ for an admissible relaxed control $\pi$.
@@
The goal of the agent is to maximize $J^{\pi}$ over all admissible relaxed controls. The non-exponential discounting renders the stochastic control problem time inconsistent. As the global optimality fails, we look for the subgame perfect Nash equilibrium for the intra-personal game such that the future selves will not turn over. The game-theoretic thinking yields the following definition.
@@
$\limsup_{\varepsilon\to 0+}\frac{J^{\pi^{\prime}\otimes_{\varepsilon}\pi^{*}}(x)-J^{\pi^{*}}(x)}{\varepsilon}\leq 0,\quad\forall\pi^{\prime}$ (2.4)
This paper aims to study the existence of equilibrium by employing the vanishing entropy regularization approach. Such an approach is first proposed in showing the general existence of equilibrium for time-inconsistent MDP problems in [1]. In the existing literature, the existence of equilibrium is typically addressed by resorting to the existence of classical solution to the extended HJB or EHJB and some rigorous verification arguments, see Proposition 2.1 below, whose proof is omitted because it is analogous to that of Proposition 3.1 later.
@@
Let Assumptions 2.1 and 2.2 hold. Suppose that there exists a classic solution $u(t,x)\in{\mathcal{C}}^{1,2}(\mathbb{T}\times\mathbb{R}^{d})$ to the nonlinear and nonlocal PDE system
$+\sup_{\varpi\in{\mathcal{P}}(U)}\left\{\int_{U}\left[b(x,a)D_{x}u(0,x)+r(0,x,a)\right]\varpi(da)\right\},$ (2.5)
and
@@
where the relaxed control $\pi^{*}:\mathbb{R}^{d}\mapsto{\mathcal{P}}(U)$ achieves the supremum of the equation (2.5). We have that $\pi^{*}$ is an equilibrium and $V^{\pi^{*}}=u$, and we call (2.5)-(2.6) the equilibrium HJB (EHJB) system.
@@
It is well known that showing the existence of a classical solution to the nonlinear and nonlocal EHJB system (2.5)–(2.6) is extremely challenging. In the existing literature, classical solutions have only been constructed in some specific concrete models, typically via a guess-and-verify approach.
In contrast, one of our main findings is to show that searching for a classical solution may not be needed. In fact, the vanishing entropy regularization approach gives the existence of a weak solution to a generalized EHJB equation, for which we can still exercise the verification arguments, thereby confirming the existence of equilibrium; see Theorem 4.1 and Corollary 4.1.
3 Time-inconsistent Control under Entropy Regularization
@@
where $\mathcal{H}$ stands for the Shannon entropy: $\mathcal{H}(\varpi):=-\int_{U}\ln(\varpi(a))\varpi(a)da$ for a given density $\varpi\in{\mathcal{P}}_{c}(U)$. For technical convenience, we shall impose some integrability assumptions on the general discounting function $\delta(t)$ in the following sense.
@@
$K_{5}:=\int_{0}^{\infty}\sup_{u\geq s}\frac{|\delta_{t}(s)-\delta_{t}(u)|}{|s-u|^{\alpha/2}}ds<\infty.$ (3.3)
Note that (2.1) in Assumption 2.1, together with (3.2), ensures that $V^{\pi}_{\lambda}$ is well-defined for all admissible policies $\pi$.
@@
Proposition 3.1.
If a feedback relaxed control $\pi^{*}:\mathbb{R}^{d}\mapsto{\mathcal{P}}(U)$ is a regularized equilibrium with entropy parameter $\lambda$ and $V^{\pi^{*}}_{\lambda}\in{\mathcal{C}}^{1,2}_{\alpha/2,\alpha}(\mathbb{T}\times\mathbb{R}^{d})$, then $V^{\pi^{*}}$ solves the following exploratory equilibrum HJB (EEHJB) system
@@
$+\sup_{\varpi\in{\mathcal{P}}_{c}(U)}\left\{\int_{U}\left[b(x,a)D_{x}u(0,x)+r(0,x,a)-\lambda\ln(\varpi(a))\right]\varpi(a)da\right\}$
$=$ $\partial_{t}u_{t}(0,x)+\frac{1}{2}\operatorname{tr}(\sigma\sigma^{T}(x)D^{2}_{x}u(0,x)+\lambda\ln\left\{\int_{U}\exp\left(\frac{1}{\lambda}\left[b(x,a)D_{x}u(0,x)+r(0,x,a)\right]\right)\right\}$ (3.5)
$0=$ $\partial_{t}u(t,x)+\frac{1}{2}\operatorname{tr}(\sigma\sigma^{T}(x)D^{2}_{x}u(t,x)+b^{\pi^{*}}(x)D_{x}u(t,x)+r^{\pi^{*}}(t,x)+\lambda\mathcal{H}(\pi^{*}(x))$ (3.6)
@@
$\pi^{*}(x,a)=\frac{\exp\left(\frac{1}{\lambda}\left[b(x,a)\cdot D_{x}u(0,x)+r(0,x,a)\right]\right)}{\int_{U}\exp\left(\frac{1}{\lambda}\left[b(x,a^{\prime})\cdot D_{x}u(0,x)+r(0,x,a^{\prime})\right]\right)da^{\prime}},\quad\forall x\in\mathbb{R}.$ (3.7)
Conversely, if a feedback relaxed control $\pi^{*}$ achieves the supremum in (3.5) and a function $u(t,x)\in{\mathcal{C}}^{1,2}_{\alpha/2,\alpha}(\mathbb{T}\times\mathbb{R}^{d})$ satisfies the EEHJB system (3.5)–(3.6), then $\pi^{*}$ is an equilibrium with entropy weight $\lambda$ satisfying (3.7) and $u(t,x)=V^{\pi^{*}}_{\lambda}(t,x)$.
Proof.
Part (i): Suppose $\pi^{*}:\mathbb{R}^{d}\mapsto{\mathcal{P}}(U)$ is a regularized equilibrium such that $V^{\pi^{*}}_{\lambda}\in{\mathcal{C}}^{1,2}_{\alpha/2,\alpha}(\mathbb{T}\times\mathbb{R}^{d})$, we will show that (3.5)–(3.7) hold.
@@
$E_{x}\left[\int_{0}^{\varepsilon}\left[r^{\pi^{*}}(t+s,X^{\pi^{*}}_{s})+\delta(t+s)\lambda\mathcal{H}(\pi^{*}(X^{\pi^{*}}_{s}))\right]ds+V^{\pi^{*}}_{\lambda}(t+\varepsilon,X^{\pi^{*}}_{s})\right]-V^{\pi^{*}}_{\lambda}(t,x)=0.$
In light that $V^{\pi^{*}}_{\lambda}\in{\mathcal{C}}^{1,2}_{\alpha/2,\alpha}(\mathbb{T}\times\mathbb{R}^{d})$, the classic Itô formula yields
@@
$J_{\lambda}^{\varpi\otimes_{\varepsilon}\pi^{*}}(x)=V^{\varpi\otimes_{\varepsilon}\pi^{*}}_{\lambda}(0,x)=\mathbb{E}\left[\int_{0}^{\varepsilon}\left[r^{\varpi}(s,X^{\varpi}_{s})+\delta(t+s)\lambda\mathcal{H}(\varpi))\right]ds+V^{\pi^{*}}_{\lambda}\left(t+\varepsilon,X^{\varpi}_{s}\right)\right].$
@@
$\frac{J^{\varpi\otimes_{\varepsilon}\pi^{*}}_{\lambda}(x)-J^{\varpi\otimes_{\varepsilon}\pi^{*}}_{\lambda}(x)}{\varepsilon}=$ $\frac{E_{x}\left[\int_{0}^{\varepsilon}\left[r^{\varpi}(s,X^{\varpi}_{s})+\delta(s)\lambda\mathcal{H}(\varpi)\right]ds\right]}{\varepsilon}$ (3.9)
$+\frac{\mathbb{E}_{x}\left[V^{\pi^{*}}_{\lambda}(\varepsilon,X^{\varpi}_{s})\right]-V^{\pi^{*}}_{\lambda}(0,x)}{\varepsilon},$
Noting that $V^{\pi^{*}}_{\lambda}\in{\mathcal{C}}^{1,2}_{\alpha/2,\alpha}(\mathbb{T}\times\mathbb{R}^{d})$, as $\varepsilon\to 0$, we deduce that the right-hand side tends to
$\partial_{t}V^{\pi^{*}}_{\lambda}(0,x)+\frac{1}{2}\operatorname{tr}\left(\sigma\sigma^{T}(x)D_{x}^{2}V^{\pi^{*}}_{\lambda}(0,x)\right)+\int_{U}\left[b(x,a)D_{x}V^{\pi^{*}}_{\lambda}(0,x)+r(0,x,a)-\lambda\ln(\varpi(a))\right]\varpi(d)da.$ (3.10)
As $\pi^{*}$ is a regularized equilibrium, we conclude that
@@
Part (ii): We now show the converse direction. Suppose that $\pi^{*}:\mathbb{R}^{d}\mapsto{\mathcal{P}}(U)$ is measurable and $u(t,x)$ $\in$ ${\mathcal{C}}^{1,2}_{\alpha/2,\alpha}(\mathbb{T}\times\mathbb{R}^{d})$ satisfies the EEHJB system (3.5)—(3.6). We first show that $V^{\pi^{*}}_{\lambda}(t,x)=u(t,x)$ for any $(t,x)\in\mathbb{T}\times\mathbb{R}^{d}$. Leveraging the condition $b^{\pi^{*}}\in L^{\infty}({\mathbb{R}^{d}})$ and the conditions on $\sigma$ given in Assumption 2.2, we deduce that the SDE
$dX^{\pi^{*}}_{t}=b^{\pi^{*}}(X_{t}^{\pi^{*}})dt+\sigma(X^{\pi^{*}}_{t})dW_{t}\quad\text{with }X_{0}=x\in{\mathbb{R}^{d}}$
admits a unique strong solution(see [26, Theorem 1]) with a density function for any positive time (see, e.g., [23, Theorem 9.1.9]). Then, for each $(t,x)\in\mathbb{T}\times{\mathbb{R}^{d}}$, applying the generalized Itô-Krylov formula (see e.g. [16, Section 2.10, Theorem 1] in a more general setting) to $u(t+s,X^{\pi^{*}}_{s})$ yields
$du(t+s,X^{\pi^{*}}_{s})$
$=\left(\partial_{t}u(t+s,X^{\pi^{*}}_{s})+b^{\pi^{*}}(x)D_{x}u(t+s,X^{\pi^{*}}_{s})+\frac{1}{2}\operatorname{tr}\bigg((\sigma\sigma^{T})(X^{\pi^{*}}_{s})D^{2}_{x}u(t+s,X^{\pi^{*}}_{s})\bigg)\right)ds$
$\qquad+\sigma(X^{\pi^{*}}_{s})D_{x}u(X^{\pi^{*}}_{s})dW_{s}.$
Taking expectations on both sides leads to
$u(t,x)=$ $\mathbb{E}_{x}\bigg[\int_{0}^{T}\left(u_{t}(t+s,X^{\pi^{*}}_{s})+b^{\pi^{*}}(x)D_{x}u(t+s,X^{\pi^{*}}_{s})+\frac{1}{2}\operatorname{tr}\left((\sigma\sigma^{T})(X^{\pi^{*}}_{s})D^{2}_{x}u(t+s,X^{\pi^{*}}_{s})\right)\right)ds$
$\qquad+u(t+T,X^{\pi^{\infty}}_{T})\bigg]$
$=$ $\mathbb{E}_{x}\left[\int_{0}^{T}r^{\pi^{*}}(t+s,X^{\pi^{*}}_{s})ds\right]+\mathbb{E}_{x}\left[u(t+T,X^{\pi^{\infty}}_{T})\right]$
$\overset{T\to\infty}{\longrightarrow}$ $\mathbb{E}_{x}\left[\int_{0}^{\infty}r^{\pi^{*}}(t+s,X^{\pi^{*}}_{s})ds\right]=V^{\pi^{*}}_{\lambda}(t,x),$
[…]
To address the general existence of classical solution to the EEHJB system (3.5)—(3.6), we need one more assumption that provides regularity of parameters on action variable $a$ and the action space $U$. To properly state the assumption, we illustrate the following notation for a cone in $\mathbb{R}^{\ell}$. When $\ell>1$, for any $\gamma\in[0,\pi/2]$, denote
Assumption 3.1.
The functions $a\mapsto b(x,a),r(t,x,a)$ are uniformly Lipchitz, i.e.,
Remark 3.1.
We also want to emphasize that all our analysis work, hence, our results hold for $U$ being finite or countable.
Theorem 3.1.
Let Assumptions 2.1, 2.2 and 3.1 hold. Then there exists a $\lambda_{0}>0$ such that for all entropy parameter $0<\lambda\leq\lambda_{0}$, a regularized equilibrium $\pi^{*}$ with entropy parameter $\lambda$ exists such that $V^{\pi^{*}}\in{\mathcal{C}}^{1,2}_{\alpha/2,\alpha}(\mathbb{T}\times\mathbb{R}^{d})$ and
$\|V^{\pi^{*}}_{\lambda}\|_{{\mathcal{C}}^{0,1}_{\alpha/2,\alpha}(\mathbb{T}\times\mathbb{R}^{d})}\vee\|V^{\pi^{*}}_{\lambda}\|_{W^{1,2,\operatorname{\textbf{ul}}}_{p}(\mathbb{T}\times\mathbb{R}^{d})}\vee\sup_{x\in\mathbb{R}^{d}}\|V^{\pi^{*}}_{\lambda}(\cdot,x)\|_{{\mathcal{C}}^{1}_{\alpha/2}(\mathbb{T})}\leq A^{*};$ (3.11)
$\|V^{\pi^{*}}_{\lambda}\|_{{\mathcal{C}}^{1,2}_{\alpha/2,\alpha}(\mathbb{T}\times\mathbb{R}^{d})}\leq A_{\lambda}.$
where $A^{*}$ is a constant depending on $K_{0},\cdots,K_{5}$, $\eta$, $\Theta,\zeta,\gamma$ in Assumptions 2.1, 2.2, 3.1, and conditions (3.2)–(3.3), but independent of $\lambda$, and $A_{\lambda}$ is another constant that also depends on $\lambda$.
3.1 Proof for Theorem 3.1
As stated in Proposition 3.1, a regularized equilibrium with the entropy parameter $\lambda>0$ is a density functional on $D_{x}V^{\pi^{*}}(0,x)$ in Gibbs-form. Take any $\lambda>0$, define $\Gamma_{\lambda}:\mathbb{R}^{d}\times\mathbb{R}^{d}\mapsto{\mathcal{P}}_{c}(U)$ by
$\Gamma_{\lambda}(x,y,a)$= $\operatorname*{arg\,max}_{\varpi\in{\mathcal{P}}_{c}(U)}\left\{\int_{U}\left[b(x,a)y+r(0,x,a)-\lambda\ln(\varpi(a))\right]\varpi(a)da\right\}$ (3.12)
$=$ $\frac{\exp\left(\frac{1}{\lambda}\left[b(x,a)y+r(0,x,a)\right]\right)}{\int_{U}\exp\left(\frac{1}{\lambda}\left[b(x,a^{\prime})y+r(0,x,a^{\prime})\right]\right)da^{\prime}}.$
Then (3.7) gives that $\pi^{*}(x,a)=\Gamma_{\lambda}(x,D_{x}V^{\pi^{*}}(0,x),a)$ if $\pi^{*}$ is a regularized equilibrium.
The following lemma is borrowed from [1, Lemma 1] (see also [12, Corollary 4.14]), which provides a sublinear growth estimate for the entropy term on the Gibbs-form operator $\Gamma_{\lambda}(x,y,a)$ in terms of $|y|$.

### 04 Apr 2026 v2

@@ 2 Model Setup
We consider a time-inconsistent stochastic control problem. Let $U$ denote the action space, which is a compact subset of $\mathbb{R}^{\ell}$ with $\operatorname{\texttt{Leb}}(U)>0$. We also denote by ${\mathcal{P}}(U)$ as the set of all probability measures on $U$ equipped with the weak topology, and ${\mathcal{P}}_{c}(U)$ as the subset of ${\mathcal{P}}(U)$ which contains all probability measures absolutely continuous with respect to the Lebesgue measure. For technical convenience, we focus on the model over an infinite horizon where only the drift coefficient of the state process is controlled. The general case when the diffusion coefficient is also controlled will be left for future study.
@@
admits a unique strong solution, where $b:\mathbb{R}^{d}\times U\to\mathbb{R}^{d}$, and $\sigma:\mathbb{R}^{d}\to\mathbb{R}^{d\times m}$ and $W$ is a standard $m$-dimensional Brownian motion.
The value function under the relaxed control $\pi$ over an infinite horizon is defined by
$V^{\pi}(t,x)=\mathbb{E}_{x}\left[\int_{0}^{\infty}\int_{U}r(t+s,X^{\pi}_{s},a)\pi_{s}(da)ds\right],$
where $r(t,x,a):\mathbb{T}\times\mathbb{R}^{d}\times U\to\mathbb{R}$ is the reward function. Let $J^{\pi}(x):=V^{\pi}(0,x)$ be the value function at the initial time $t=0$ under the relaxed control $\pi$.
For the rest of the paper, we use notation
$f^{\varpi}(t,x):=\int_{U}f(t,x,a)\varpi(da)$
for any generic function $f(t,x,a):\mathbb{T}\times{\mathbb{R}^{d}}\times U\to\mathbb{R}$ and any generic distribution $\varpi\in{\mathcal{P}}(U)$. And we use the shorthand $f^{\pi}(x)=f^{\pi(x)}(t,x)$ for a feedback control $\pi:\mathbb{R}^{d}\to{\mathcal{P}}(U)$.
@@
$K_{2}:=\int_{0}^{\infty}\sup_{x\in\mathbb{R}^{d},a\in U,s<u\leq s+1}\frac{|\partial_{t}r(s,x,a)-\partial_{t}r(u,x,a)|}{|s-u|^{\alpha/2}}ds<\infty.$ (2.2)
@@
The boundedness condition in (2.1) ensures that $J^{\pi}$ is well-defined for any admissible relaxed control $\pi$, and the condition in (2.2) is satisfied when the second derivative of $r$ with respect to $t$ is integrable over the time horizon $\mathbb{T}$, uniformly in $a$ and $x$. This assumption will be used later to ensure the Hölder continuity of $\partial_{t}V^{\pi}$ for an admissible relaxed control $\pi$.
@@
The goal of the agent is to maximize $J^{\pi}$ over all admissible relaxed controls. The non-exponential discounting renders the stochastic control problem time inconsistent. Since global optimality is no longer attainable, we seek a subgame perfect Nash equilibrium for the intra-personal game under which future selves have no incentive to deviate from the prescribed strategy. This game-theoretic perspective motivates the following definition.
@@
$\limsup_{\varepsilon\to 0+}\frac{J^{\pi^{\prime}\otimes_{\varepsilon}\pi^{*}}(x)-J^{\pi^{*}}(x)}{\varepsilon}\leq 0,\quad\forall\text{ admissible }\pi^{\prime}.$ (2.4)
This paper aims to study the existence of equilibrium by employing the vanishing entropy regularization approach. Such an approach is first proposed in showing the general existence of equilibrium for time-inconsistent MDP problems in [1]. In the existing literature, the existence of equilibrium is typically addressed by resorting to the existence of classical solution to the extended HJB or EHJB and some rigorous verification arguments, see Proposition 2.1 below, whose proof is omitted because it is analogous to that of Theorem 3.1 Part (i).
@@
Let Assumptions 2.1 and 2.2 hold. Suppose that there exists a classic solution $u(t,x)\in{\mathcal{C}}^{1,2}(\mathbb{T}\times\mathbb{R}^{d})$ to the following nonlinear and nonlocal PDE system
$0=$ $\partial_{t}u(0,x)+\frac{1}{2}\operatorname{tr}\left(\sigma\sigma^{T}(x)D^{2}_{x}u(0,x)\right)+\sup_{\varpi\in{\mathcal{P}}(U)}\left\{\int_{U}\left[b(x,a)D_{x}u(0,x)+r(0,x,a)\right]\varpi(da)\right\},$ (2.5)
@@
where the relaxed control $\pi^{*}:\mathbb{R}^{d}\to{\mathcal{P}}(U)$ achieves the supremum of the equation (2.5). Then we have that $\pi^{*}$ is an equilibrium and $V^{\pi^{*}}=u$. We call (2.5)–(2.6) the equilibrium HJB (EHJB) system.
@@
It is well known that proving the existence of a classical solution to the nonlinear and nonlocal EHJB system (2.5)–(2.6) is extremely challenging. In the existing literature, classical solutions have only been constructed in some specific concrete models, typically via a guess-and-verify approach.
In contrast, one of our main findings is that searching for a classical solution is unnecessary. Indeed, the vanishing entropy regularization approach yields the existence of a strong solution (rather than a classical one) to the EHJB system, for which verification arguments remain applicable, thereby establishing the existence of an equilibrium; see Theorems 4.2 and 4.1.
3 Existence of Equilibrium under Entropy Regularization
@@
where $\mathcal{H}$ stands for the Shannon entropy: $\mathcal{H}(\varpi):=-\int_{U}\ln(\varpi(a))\varpi(a)da$ for a given density $\varpi\in{\mathcal{P}}_{c}(U)$. For technical convenience, we impose the following integrability assumptions on the general discounting function $\delta(t)$.
@@
$K_{5}:=\int_{0}^{\infty}\sup_{s<u\leq s+1}\frac{|\delta_{t}(s)-\delta_{t}(u)|}{|s-u|^{\alpha/2}}ds<\infty.$ (3.3)
Note that (2.1) in Assumption 2.1, together with (3.2), ensures that $V^{\pi}_{\lambda}$ is well-defined for all admissible policies $\pi\in{\mathcal{P}}_{c}(U)$.
@@
To address the general existence of classical solution to the EEHJB system (3.5)–(3.6), we need one more assumption that provides regularity of parameters on action variable $a$ and the action space $U$. To properly state the assumption, we illustrate the following notation for a cone in $\mathbb{R}^{\ell}$. When $\ell>1$, for any $\gamma\in[0,\pi/2]$, denote
Assumption 3.1.
The functions $a\mapsto b(x,a)$ and $a\mapsto r(t,x,a)$ are uniformly Lipchitz, i.e.,
Remark 3.1.
Theorem 3.1.
Assume Assumptions 2.1, 2.2, 3.1, (3.2) and (3.3) hold. Then the followings hold.
  • (i)
    If a feedback relaxed control $\pi^{*}:\mathbb{R}^{d}\to{\mathcal{P}}_{c}(U)$ is a regularized equilibrium with entropy parameter $\lambda$ and $V^{\pi^{*}}_{\lambda}\in{\mathcal{C}}^{1,2}_{\alpha/2,\alpha}(\mathbb{T}\times\mathbb{R}^{d})$, then $V^{\pi^{*}}$ solves the following exploratory equilibrium HJB (EEHJB) system
@@
    $+\sup_{\varpi\in{\mathcal{P}}_{c}(U)}\left\{\int_{U}\left[b(x,a)D_{x}u(0,x)+r(0,x,a)-\lambda\ln(\varpi(a))\right]\varpi(a)da\right\}$ (3.5)
    $=$ $\partial_{t}u_{t}(0,x)+\frac{1}{2}\operatorname{tr}(\sigma\sigma^{T}(x)D^{2}_{x}u(0,x)$
    $+\lambda\ln\left\{\int_{U}\exp\left(\frac{1}{\lambda}\left[b(x,a)D_{x}u(0,x)+r(0,x,a)\right]\right)\right\},\forall x\in\mathbb{R}^{d};$
    $0=$ $\partial_{t}u(t,x)+\frac{1}{2}\operatorname{tr}(\sigma\sigma^{T}(x)D^{2}_{x}u(t,x))$
    $+b^{\pi^{*}}(x)D_{x}u(t,x)+r^{\pi^{*}}(t,x)+\lambda\delta(t)\mathcal{H}(\pi^{*}(x)),\forall(t,x)\in\mathbb{T}\times\mathbb{R}^{d},$ (3.6)
@@
    $\pi^{*}(x,a)=\frac{\exp\left(\frac{1}{\lambda}\left[b(x,a)\cdot D_{x}u(0,x)+r(0,x,a)\right]\right)}{\int_{U}\exp\left(\frac{1}{\lambda}\left[b(x,a^{\prime})\cdot D_{x}u(0,x)+r(0,x,a^{\prime})\right]\right)da^{\prime}}\quad\forall x\in\mathbb{R},a\in U.$ (3.7)
  • (ii)
    Conversely, if a couple $(u,\pi^{*})$ satisfies that $u(t,x)\in{\mathcal{C}}^{1,2}_{\alpha/2,\alpha}(\mathbb{T}\times\mathbb{R}^{d})$ being a classic solution to the EEHJB system (3.5)–(3.6) and (3.7) holds. Then, $\pi^{*}$ is a regularized equilibrium with entropy weight $\lambda$ and $u(t,x)=V^{\pi^{*}}_{\lambda}(t,x)$.
  • (iii)
    There exists $\lambda_{0}>0$ such that for all entropy parameters $0<\lambda\leq\lambda_{0}$, there exists a pair $(u,\pi^{*})$ with $u(t,x)\in{\mathcal{C}}^{1,2}_{\alpha/2,\alpha}(\mathbb{T}\times\mathbb{R}^{d})$ solving the EEHJB system (3.5)–(3.6), and $\pi^{*}$ satisfies (3.7). As a consequence, $u(t,x)=V^{\pi^{*}}_{\lambda}(t,x)$, and $\pi^{*}$ is an equilibrium under entropy weight $\lambda$. Moreover, the following estimates also hold:
    $\|V^{\pi^{*}}_{\lambda}\|_{{\mathcal{C}}^{0,1}_{\alpha/2,\alpha}([t,\infty)\times\mathbb{R}^{d})}\vee\|V^{\pi^{*}}_{\lambda}\|_{W^{1,2,\operatorname{\textbf{ul}}}_{p}([t,\infty)\times\mathbb{R}^{d})}\leq A^{*}\Psi(t)\quad\forall t\in\mathbb{T};$ (3.8)
    $\sup_{x\in\mathbb{R}^{d}}\|V^{\pi^{*}}_{\lambda}(\cdot,x)\|_{{\mathcal{C}}^{1}_{\alpha/2}(\mathbb{T})}\vee\|V^{\pi^{*}}_{\lambda}\|_{{\mathcal{C}}^{1,2}_{\alpha/2,\alpha}(\mathbb{T}\times\mathbb{R}^{d})}\leq A_{\lambda}.$
    Here $A^{*}$ is a constant depending on $K_{0},\cdots,K_{5}$, $\eta$, $\Theta,\zeta,\gamma$ in Assumptions 2.1, 2.2, 3.1, and conditions (3.2)–(3.3), but independent of $\lambda$, and $A_{\lambda}$ is another constant that also depends on $\lambda$. $\Psi$ is a deterministic function from $\mathbb{T}$ to $\mathbb{T}$ that satisfies $\Psi(0)=1$, $\Psi$ decreases and $\lim_{t\to\infty}\Psi(t)=0$.
3.1 Proof for Theorem 3.1
We begin by proving Part (i) of Theorem 3.1. The proofs of Parts (ii) and (iii) will be presented after establishing several auxiliary lemmas. In particular, the proof of Part (ii) will follow Lemma 3.3, while the proof of Part (iii) will be delivered after Lemma 3.5.
Proof for Theorem 3.1 Part (i): .
Suppose $\pi^{*}:\mathbb{R}^{d}\to{\mathcal{P}}_{c}(U)$ is a regularized equilibrium such that $V^{\pi^{*}}_{\lambda}\in{\mathcal{C}}^{1,2}_{\alpha/2,\alpha}(\mathbb{T}\times\mathbb{R}^{d})$. We will show that (3.5)–(3.7) hold.
@@
$E_{x}\left[\int_{0}^{\varepsilon}\left[r^{\pi^{*}}(t+s,X^{\pi^{*}}_{s})+\delta(t+s)\lambda\mathcal{H}(\pi^{*}(X^{\pi^{*}}_{s}))\right]ds+V^{\pi^{*}}_{\lambda}(t+\varepsilon,X^{\pi^{*}}_{\varepsilon})\right]-V^{\pi^{*}}_{\lambda}(t,x)=0.$
In light that $V^{\pi^{*}}_{\lambda}\in{\mathcal{C}}^{1,2}_{\alpha/2,\alpha}(\mathbb{T}\times\mathbb{R}^{d})$, the classical Itô formula yields
@@
$J_{\lambda}^{\varpi\otimes_{\varepsilon}\pi^{*}}(x)=V^{\varpi\otimes_{\varepsilon}\pi^{*}}_{\lambda}(0,x)=\mathbb{E}\left[\int_{0}^{\varepsilon}\left[r^{\varpi}(s,X^{\varpi}_{s})+\delta(s)\lambda\mathcal{H}(\varpi))\right]ds+V^{\pi^{*}}_{\lambda}\left(\varepsilon,X^{\varpi}_{\varepsilon}\right)\right].$
@@
$\frac{J^{\varpi\otimes_{\varepsilon}\pi^{*}}_{\lambda}(x)-J^{\pi^{*}}_{\lambda}(x)}{\varepsilon}=$ (3.10)
$\frac{1}{\varepsilon}\left\{\mathbb{E}_{x}\left[\int_{0}^{\varepsilon}\left(r^{\varpi}(s,X^{\varpi}_{s})+\delta(s)\lambda\mathcal{H}(\varpi)\right)ds\right]+\mathbb{E}_{x}\left[V^{\pi^{*}}_{\lambda}(\varepsilon,X^{\varpi}_{\varepsilon})\right]-V^{\pi^{*}}_{\lambda}(0,x)\right\}.$
Again applying Itô formula to $V^{\pi^{*}}_{\lambda}\in{\mathcal{C}}^{1,2}_{\alpha/2,\alpha}(\mathbb{T}\times\mathbb{R}^{d})$, we obtain that the limit of the right-hand side as $\varepsilon\to 0$ is
$\partial_{t}V^{\pi^{*}}_{\lambda}(0,x)+\frac{1}{2}\operatorname{tr}\left(\sigma\sigma^{T}(x)D_{x}^{2}V^{\pi^{*}}_{\lambda}(0,x)\right)$ (3.11)
$+\int_{U}\left[b(x,a)D_{x}V^{\pi^{*}}_{\lambda}(0,x)+r(0,x,a)-\lambda\ln(\varpi(a))\right]\varpi(d)da.$
By the definition of regularized equilibrium, we conclude that
@@
As stated in Theorem 3.1 Part (i), a regularized equilibrium with the entropy parameter $\lambda>0$ is a density functional on $D_{x}V^{\pi^{*}}(0,x)$ in Gibbs form. For any fixed $\lambda>0$, define $\Gamma_{\lambda}:\mathbb{R}^{d}\times\mathbb{R}^{d}\to{\mathcal{P}}_{c}(U)$ by
$\Gamma_{\lambda}(x,p,a)$= $\operatorname*{arg\,max}_{\varpi\in{\mathcal{P}}_{c}(U)}\left\{\int_{U}\left[b(x,a)p+r(0,x,a)-\lambda\ln(\varpi(a))\right]\varpi(a)da\right\}$ (3.12)
$=$ $\frac{\exp\left(\frac{1}{\lambda}\left[b(x,a)p+r(0,x,a)\right]\right)}{\int_{U}\exp\left(\frac{1}{\lambda}\left[b(x,a^{\prime})p+r(0,x,a^{\prime})\right]\right)da^{\prime}}.$
The following lemma is borrowed from [1, Lemma 1] (also see [12, Corollary 4.14]), which provides a sublinear growth estimate for the entropy term on the Gibbs-form operator $\Gamma_{\lambda}(x,p,a)$ in terms of $|p|$.